Unit and dimension
1
Expert's answer
2021-09-27T18:28:15-0400

Dimensions:-

A dimension is a measure of a physical variable (without numerical values).

For example

Mass , volume

Unit:- a unit is a way to assign a number or measurement to that dimension.

For example


length is a dimension, but it is measured in units of feet (ft) or meters (m).

Magnetic flux "\phi=B.A"

Where B= magnetic field

A=area

Unit Weber

Magnetic field (B)=Wb/m2


Another example

Electric field

"E=\frac{V}{l}"

Unit "=\frac{Volt}{meter}"